Solve by using the quadratic formula.
3x² = 11x + 4
3x^2-11x-4 11+/- squareroot (-11)^2-4(3)(-4)/2(3)
11+/- squareroot 121+48/6
11+/- 13/6
24/6= 4 or -2/6= -1/3

Answer by Difficulty(2)

3x^2-11x-4 11+/- squareroot (-11)^2-4(3)(-4)/2(3)
11+/- squareroot 121+48/6
11+/- 13/6
24/6= 4 or -2/6= -1/3